The intensity of a wave at a point is 5.0 W m -2 when the amplitude of the wave there is 2.0 mm. What is

the amplitude at the point when the intensity is increased to 22.5 W m -2 ?


1
Expert's answer
2021-10-24T18:23:59-0400

The intensity of a sound is proportional the amplitude squared. Hence:


"\frac{I_1}{I_2}=\frac{A_1^2}{A_2^2},\\\space\\
A_2=A_1\sqrt\frac{I_2}{I_1}=4.24\text{ mm}."